# references.bib

### v3

@@ 3.3 Free Heyting Algebras from Distributive Lattices
Given a $g$-Vietoris complex over $X$, one can form the projective limit of the sequence, which, since Priestley spaces are closed under such a construction, will again be a Priestley space. We denote this limit by $V_{G}^{g}(X)$ (the $G$ standing for “Ghilardi”). The purpose of such a construction lies in the universal property which is carried out over any one step, which we proceed to outline:
Lemma 17.
@@
Lemma 18.
The property of Lemma 17 is equivalent to the following: given a Priestley space $Z$ with a $g$-open continuous and order-preserving map $h:Z\to X$, there exists a unique $r_{g}$-open, continuous and order-preserving map $h^{\prime}$ such that the triangle in Figure 4 commutes.
@@
This follows immediately from Lemma 12 by dualising the statements. ∎
@@
(Of Lemma 17) We show the property from Lemma 18. Given $Z,X$ Priestley spaces as given, first we note that the definition of $h^{\prime}$ is totally forced by the commutativity of the triangle: if $k:Z\to V_{g}(X)$ is any map in these conditions, we will show that
@@
Lemma 19.
Let $p:X\to Y$ be a monotone and continuous function between Priestley spaces, and $g_{X}:X\to Z$ and $g_{Y}:Y\to Z$ be two maps such that the obvious triangle commutes, and the relative pseudocomplements indexed by $g_{X}$ and $g_{Y}$ exist. Then $p[-]:V_{g}(X)\to V_{g}(Y)$ is a unique monotone and continuous function such that for each $C\in V_{g_{x}}(X)$, $pr(C)=r_{g_{Y}}p[C]$.
@@
It is clear that $p[-]$ will be a monotone and continuous function. To see its uniqueness, note that $p[-]$ is $r_{g_{Y}}$-open, and hence by Lemma 17, it is the unique extension of $pr$ (which is $g_{Y}$-open given that $p$ is $g_{Y}$-open). ∎
@@
Using repeatedly Lemma 17, given that $Z$ is an Esakia space, we get a sequence of maps, starting with $h_{0}=h$, and letting $h_{n+1}=h_{n}[-]$. Then for each $x\in Z$, we have that $\overline{h}(x)$ will be the element
@@
We use Lemma 19 repeatedly on the $g_{X}$-Vietoris complex over $X$ and the $g_{Y}$-Vietoris complex over $Y$: we let $p_{0}=p$, and $p_{n+1}=p[-]$. This lifts to a function on the inverse limits mapping $x\in V_{G}(X)$ to
@@
It is not difficult to see that having $\overline{p}$ so defined, $\pi_{0}\overline{p}$ is a $g_{Y}$-open map, so $\overline{p}$ will additionally coincide with the unique p-morphism lifting of $p\pi_{0}$ (by Lemma 20), in addition to making all diagrams commute. ∎
@@
Theorem 22.
The functor $V_{G}$ is right adjoint to the inclusion $I:\mathbf{Esa}\to\mathbf{Pries}$. Moreover,
@@
This follows immediately from Lemma 20 and Lemma 21. Note that the unit $\eta_{X}:X\to V_{G}(X)$ consists of the unique lifting of the identity $id_{X}$, and is an embedding in the category of Esakia spaces; the co-unit $\varepsilon_{X}:V_{G}(X)\to X$ is the projection onto the first coordinate, and is surjective as noted before. Hence the right adjoint is full whilst the left adjoint is faithful. ∎
@@
i.e., the free distributive lattice on the pairs $a,b\in D_{n}$, quotiented by the axioms forcing these pairs to act like relative pseudocomplements, and forcing the relative pseudocomplements added at the stage $n+2$ to agree with those added at stage $n+1$. We can then consider the homomorphism $i_{n}:D_{n}\to D_{n+1}$ which sends $a$ to the equivalence class $[(1,a)]_{\Theta}$. This gives us a directed system $(D_{n},i_{n})$ which, as is easy to see, has the same universal properties as the Vietoris complex over $X$, where $X$ is the dual Priestley space of $D$, and in fact this means that $i_{n}$ are all inclusions. By Lemma 8, we have that forming the directed limit is dual to taking the inverse limit, i.e., $\lim_{n\in\omega}D_{n}$ is dual to $V_{G}(X)$. Let $F_{G}(D)=\lim_{n\in\omega}D_{n}$ as stated. Moreover, $F_{G}$ forms a functor as well, obtained by dualizing all the above facts. Hence we have:
Theorem 23.
@@
The topology will again be given by the finite subsets not containing $\omega$, and the cofinite subsets containing it. Proceeding in this way, we obtain spaces $V_{2}(X)$, $V_{3}(X)$, etc, such that each of them is topologically isomorphic to the Alexandroff compactification of $\omega$, and order theoretically, on each root, it looks like the $n$-th step construction of the Rieger-Nishimura. Then $V_{G}(X)$ will be isomorphic to infinitely many copies of the Rieger-Nishimura, where one of the points is identified in them, and where they are compactified by that unique point.
@@ 4.1 Free Heyting Algebras
As a first application of Theorem 23, we show how to obtain a description of the free Heyting algebra on any number of generators, and study some of its basic properties. Let $X$ be an arbitrary set, and recall Definition 6, i.e., $\mathbf{D}(X)$ being the free distributive lattice on $X$ many generators.
Theorem 25.
@@
This construction restricts, in the case of $X$ being finite, to the Ghilardi/Urquhart construction [ghilardifreeheyting, Urquhart1973]. We note that unlike that situation, it is not immediately clear that $F_{G}(\mathbf{D}(X))$ carries the additional strucutre of a bi-Heyting algebra; that is because, to show it, it would be necessary to show that $V_{1}(2^{X})$ is likewise a bi-Esakia space. Such a study involves developing substantially more theory concerning the Vietoris complex structure, and so we delay such an analysis to the sequel to this paper [openmapsandinterpolantcomplexity].
@@
By Lemma 28, $\max(V_{G}(\mathbf{2}^{X}))=\mathbf{2}^{X}$, i.e., the maximum is homeomorphic to the generalized Cantor space on $X$ many elements. If $V_{G}(\mathbf{2}^{X})$ was extremally order-disconnected, by Lemma 27 we would have that $\mathbf{2}^{X}$ would be extremally disconnected, which is not true: dually, the free Boolean algebras on $X$ generators would be complete Boolean algebras, which is known to not hold [Koppelberg1989-lg, pp.133]. ∎
@@ 4.2 Coproducts and Colimits of Heyting Algebras
consisting of the rooted, closed $\pi_{X}$-open and $\pi_{Y}$-open subsets of $X\times Y$. It follows from Lemmas 15 and 15 that $V_{1}^{\pi_{X},\pi_{Y}}(X\times Y)$ is a Priestley space, and that the root map is continuous, monotone and $\pi_{X},\pi_{Y}$-open.
We will then define the $(\pi_{X},\pi_{Y})$-Vietoris complex over $X\times Y$ in an analogous way to Definition 16, with the only difference being that we have two maps at the base. Let $V_{G}^{\pi_{X},\pi_{Y}}(X\times Y)$ be the inverse limit of that inverse system. Then we have the following:
Lemma 31.
@@
Given any Esakia space $Z$, if $f:Z\to X$ and $g:Z\to Y$ are two p-morphisms, then $(f\times g):Z\to X\times Y$ is a $\pi_{X},\pi_{Y}$-open map, and $\pi_{X}(f\times g)=f$ and $\pi_{Y}(f\times g)=g$. Using the universal property of $V_{n}^{\pi_{X},\pi_{Y}}(X\times Y)$, we obtain a unique map to each such space, such that $(f\times g)_{0}=(f\times g)$ and $(f\times g)_{n+1}=(f\times g)_{n}[-]$; this in turn forces, by Lemma 18 and the properties of inverse limits, the existence of a unique p-morphism to $(f\times g)_{\infty}:Z\to V_{G}^{\pi_{X},\pi_{Y}}(X\times Y)$, commuting with this diagram, i.e., such that $\pi_{0}(f\times g)_{\infty}=(f\times g)$. Then $p_{X}(f\times g)_{\infty}=f$ and $p_{Y}(f\times g)_{\infty}=g$, as desired. ∎
@@
For $n=0$ this is the aforementioned fact. Now assume that it holds for $n$. Given $C\in V_{n+1}^{\pi_{X},\pi_{Y\sqcup Z}}(X\times Y\sqcup Z)$, note that $C$ is a subset of $V_{n}^{\pi_{X},\pi_{Y\sqcup Z}}(X\times Y\sqcup Z)$; the latter is by induction hypothesis isomorphic to $V_{n}^{\pi_{X},\pi_{Y}}(X\times Y)\sqcup V_{n}^{\pi_{X},\pi_{Z}}(X\times Z)$, so via the isomorphism, $C$ is a subset of the disjoint union; but since $C$ is rooted, it is a subset of one of the two disjoints. This assignment is easily seen to be bijective, continuous, and monotone in both directions, giving us an isomorphism.
@@
Using similar techniques one can show that coproducts of Heyting algebras are disjoint, and that the category of Heyting algebras form an extensive category. We leave a detailed study of the full properties of such a category to further work. Instead we turn to a question of amalgamation: indeed, making use of the explicit description of the product given before, we can likewise provide, dually, a description of the pullback of two Esakia spaces:
Theorem 35.
@@
The proof of this follows exactly the same ideas as that of Theorem 32. Using this pullback we can obtain an explicit proof of the amalgamation property for Heyting algebras, which is “constructive” in the sense that it is brought about by categorical, rather than model-theoretic or logical, considerations.
Theorem 36.
@@ 4.3 Esakia Duality from Priestley Duality
As a final application we show how Esakia duality for Heyting algebras can be presented coalgebraically on the basis of Priestley by using the previous functor. For that purpose, given a Priestley space $X$, recall the map
@@
the unit of the adjunction, which is unique p-morphism lifting the identity. Explicitly, $\eta$ maps $x\in X$ to the tuple
@@
    $X$ admits a $V_{G}$-coalgebra structure.
@@
(1) implies (2): if $X$ is an Esakia space, then it admits a free $V_{G}$-coalgebra structure, namely the unit map $\eta:X\to V_{G}(X)$.

### v4

@@ 3.3 Free Heyting Algebras from Distributive Lattices
[…]
Lemma 3.8.
@@
Lemma 3.9.
The property of Lemma 3.8 is equivalent to the following: given a Priestley space $Z$ with a $g$-open continuous and order-preserving map $h:Z\to X$, there exists a unique $r_{g}$-open, continuous and order-preserving map $h^{\prime}$ such that the triangle in Figure 4 commutes.
@@
This follows immediately from Lemma 3.3 by dualising the statements. ∎
@@
(Of Lemma 3.8) We show the property from Lemma 3.9. Given $Z$, $X$ and $h$ as given, first we note that the definition of $h^{\prime}$ is totally forced by the commutativity of the triangle: if $k:Z\to V_{g}(X)$ is any map in these conditions, we will show that
@@
Lemma 3.10.
Let $p:X\to Y$ be a monotone and continuous function between Priestley spaces, and $g_{X}:X\to Z$ and $g_{Y}:Y\to Z$ be two maps such that the obvious triangle commutes, and ${\downarrow}h^{-1}[U-V]$ is clopen whenever $U,V$ are clopen upsets and $h\in\{g_{X},g_{Y}\}$. Then $p[-]:V_{g}(X)\to V_{g}(Y)$ is a unique monotone and continuous function such that for each $C\in V_{g_{x}}(X)$, $pr(C)=r_{g_{Y}}p[C]$.
@@
It is clear that $p[-]$ will be a monotone and continuous function. To see its uniqueness, note that $p[-]$ is $r_{g_{Y}}$-open, and hence by Lemma 3.8, it is the unique extension of $pr$ (which is $g_{Y}$-open given that $p$ is $g_{Y}$-open). ∎
@@
Using repeatedly Lemma 3.8, given that $Z$ is an Esakia space, we get a sequence of maps, starting with $h_{0}=h$, and letting $h_{n+1}=h_{n}[-]$. Then for each $x\in Z$, we have that $\overline{h}(x)$ will be the element
@@
We use Lemma 3.10 repeatedly on the $g_{X}$-Vietoris complex over $X$ and the $g_{Y}$-Vietoris complex over $Y$: we let $p_{0}=p$, and $p_{n+1}=p[-]$. This lifts to a function on the inverse limits mapping $x\in V_{G}(X)$ to
@@
It is not difficult to see that having $\overline{p}$ so defined, $\pi_{0}\overline{p}$ is a $g_{Y}$-open map, so $\overline{p}$ will additionally coincide with the unique p-morphism lifting of $p\pi_{0}$ (by Lemma 3.11), in addition to making all diagrams commute. ∎
@@
Theorem 3.13.
The functor $V_{G}$ is right adjoint to the inclusion $I:\mathbf{Esa}\to\mathbf{Pries}$.
@@
This follows immediately from Lemma 3.11 and Lemma 3.12. Note that the unit $\eta_{X}:X\to V_{G}(X)$ consists of the unique lifting of the identity $id_{X}$, and is an embedding in the category of Esakia spaces; the co-unit $\varepsilon_{X}:V_{G}(X)\to X$ is the projection onto the first coordinate, and is surjective as noted before. Hence the right adjoint is full whilst the left adjoint is faithful. ∎
@@
i.e., the free distributive lattice on the pairs $a,b\in D_{n}$, quotiented by the congruence forcing these pairs to act like relative pseudocomplements, and forcing the relative pseudocomplements added at the stage $n+2$ to agree with those added at stage $n+1$. We can then consider the homomorphism $i_{n}:D_{n}\to D_{n+1}$ which sends $a$ to the equivalence class $[(1,a)]_{\Theta}$. This gives us a directed system $(D_{n},i_{n})$ which, as is easy to see, has the same universal properties as the Vietoris complex over $X$, where $X$ is the dual Priestley space of $D$, and in fact this means that $i_{n}$ are all inclusions. By Lemma 2.8, we have that forming the directed limit is dual to taking the inverse limit, i.e., $\lim_{n\in\omega}D_{n}$ is dual to $V_{G}(X)$. Let $F_{G}(D)=\lim_{n\in\omega}D_{n}$ as stated. Moreover, $F_{G}$ forms a functor as well, obtained by dualizing all the above facts. Hence we have:
Theorem 3.14.
@@
The topology will again be given by the finite subsets not containing $\omega$, and the cofinite subsets containing it. Proceeding in this way, we obtain spaces $V_{2}(X)$, $V_{3}(X)$, etc, such that each of them is topologically isomorphic to the Alexandroff compactification of $\omega$, and order theoretically, on each root, it looks like the $n$-th step construction of the Rieger-Nishimura. Then $V_{G}(X)$ will be isomorphic to infinitely many copies of the Rieger-Nishimura, where the point labelled with $1$ is identified between all of them, and compactifies the disjoint union of these topologies.
@@ 4 Some Applications to the theory of Heyting Algebras
In this section we provide some basic applications of the construction of Section 3 to the theory of Heyting algebras. Whilst in subsequent sections we will turn to some novel applications, in this section we will derived known results (or results easily derived from results in the literature); however, we believe that the methodology can be suggestive to tackle similar questions in settings which may not have such a developed theory as Heyting algebras.
@@ 4.1 Free Heyting Algebras
As a first application of Theorem 3.14, we show how to obtain a description of the free Heyting algebra on any number of generators, and study some of its basic properties. Let $X$ be an arbitrary set, and recall Definition 2.6, i.e., $\mathbf{D}(X)$ being the free distributive lattice on $X$ many generators.
Theorem 4.1.
@@
This construction restricts, in the case of $X$ being finite, to the Ghilardi/Urquhart construction [ghilardifreeheyting, Urquhart1973]. We note that unlike that situation, it is not immediately clear that $F_{G}(\mathbf{D}(X))$ carries the additional structure of a bi-Heyting algebra; that is because, to show it, it would be necessary to show that $V_{1}(2^{X})$ is likewise a bi-Esakia space. Such a study involves developing substantially more theory concerning the Vietoris complex structure, and so we delay such an analysis to the sequel to this paper [openmapsandinterpolantcomplexity].
@@
By Lemma 4.4, $\max(V_{G}(\mathbf{2}^{X}))=\mathbf{2}^{X}$, i.e., the maximum is homeomorphic to the generalized Cantor space on $X$ many elements. Were $V_{G}(\mathbf{2}^{X})$ extremally order-disconnected, by Lemma 4.3 we would have that $\mathbf{2}^{X}$ would be extremally disconnected, which is not true: dually, the free Boolean algebras on $X$ generators would be complete Boolean algebras, which is known to not hold [Koppelberg1989-lg, pp.133]. ∎
@@ 4.2 Coproducts and Colimits of Heyting Algebras
consisting of the rooted, closed $\pi_{X}$-open and $\pi_{Y}$-open subsets of $X\times Y$. It follows from Lemmas 3.6 and 3.6 that $V_{1}^{\pi_{X},\pi_{Y}}(X\times Y)$ is a Priestley space, and that the root map is continuous, monotone and $\pi_{X},\pi_{Y}$-open.
We will then define the $(\pi_{X},\pi_{Y})$-Vietoris complex over $X\times Y$ in an analogous way to Definition 3.7, with the only difference being that we have two maps at the base. Let $V_{G}^{\pi_{X},\pi_{Y}}(X\times Y)$ be the inverse limit of that inverse system. Then we have the following:
Lemma 4.7.
@@
Given any Esakia space $Z$, if $f:Z\to X$ and $g:Z\to Y$ are two p-morphisms, then $(f\times g):Z\to X\times Y$ is a $\pi_{X},\pi_{Y}$-open map, and $\pi_{X}(f\times g)=f$ and $\pi_{Y}(f\times g)=g$. Using the universal property of $V_{n}^{\pi_{X},\pi_{Y}}(X\times Y)$, we obtain a unique map to each such space, such that $(f\times g)_{0}=(f\times g)$ and $(f\times g)_{n+1}=(f\times g)_{n}[-]$; this in turn forces, by Lemma 3.9 and the properties of inverse limits, the existence of a unique p-morphism to $(f\times g)_{\infty}:Z\to V_{G}^{\pi_{X},\pi_{Y}}(X\times Y)$, commuting with this diagram, i.e., such that $\pi_{0}(f\times g)_{\infty}=(f\times g)$. Then $p_{X}(f\times g)_{\infty}=f$ and $p_{Y}(f\times g)_{\infty}=g$, as desired. ∎
@@
For $n=0$ this is the aforementioned fact. Now assume that it holds for $n$. Given $C\in V_{n+1}^{\pi_{X},\pi_{Y\sqcup Z}}(X\times(Y\sqcup Z))$, note that $C$ is a subset of $V_{n}^{\pi_{X},\pi_{Y\sqcup Z}}(X\times(Y\sqcup Z))$; the latter is by induction hypothesis isomorphic to $V_{n}^{\pi_{X},\pi_{Y}}(X\times Y)\sqcup V_{n}^{\pi_{X},\pi_{Z}}(X\times Z)$, so via the isomorphism, $C$ is a subset of the disjoint union; but since $C$ is rooted, it is a subset of one of the two disjoints. This assignment is easily seen to be bijective, continuous, and monotone in both directions, giving us an isomorphism.
@@
Using similar techniques one can show that coproducts of Heyting algebras are disjoint, and that the category of Heyting algebras form a co-extensive category. This could be deduced from known results in [Carboni1993], though the explicit proof given here might be of independent interest. Finally we turn to a question of amalgamation: indeed, making use of the explicit description of the product given before, we can likewise provide, dually, a description of the pullback of two Esakia spaces:
Theorem 4.11.
@@
The proof of this follows exactly the same ideas as that of Theorem 4.8; indeed note that the projections from $V_{G}^{\pi_{X},\pi_{Y}}(X\times_{Z}Y)$ to $X$ and $Y$ are always surjective, since for each $x\in X$ (resp. $y\in Y$) ${\uparrow}x$ is a $(\pi_{X},\pi_{Y})$-open closed and rooted subset. Using this pullback we can obtain an explicit proof of the amalgamation property for Heyting algebras, which is “constructive” in the sense that it is brought about by categorical, rather than model-theoretic or logical, considerations.
Theorem 4.12.
@@ 4.3 Esakia Duality from Priestley Duality
As a final application we show how Esakia duality for Heyting algebras can be presented coalgebraically on the basis of Priestley by using the previous functor. This recovers the results from [daveycoalgebraheyting]. For that purpose, given a Priestley space $X$, recall the map
@@
the unit of the adjunction, which is the unique p-morphism lifting the identity. Explicitly, $\eta$ maps $x\in X$ to the tuple
@@
    $X$ admits a (unique) $V_{G}$-coalgebra structure.
@@
(1) implies (2): if $X$ is an Esakia space, then it admits a free $V_{G}$-coalgebra structure, namely the unit map $\eta:X\to V_{G}(X)$. Moreover, this structure is unique: if $f:X\to V_{G}(X)$ is any such coalgebra, then for each $x\in X$, $f(x)=(a_{0},a_{1},...)$; by the commutation conditions, then $\pi_{0}f(x)=x$; by the universality of the condition, then $f=\eta_{X}$.
@@
As a consequence we obtain that Esakia spaces can be seen, coalgebraically, as arising exactly from Priestley spaces by applying the $V_{G}$ construction. Moreover, we have the following:
Theorem 4.14.
The functor $V_{G}$ is comonadic.
Proof.
In light of Theorem 4.13, all that is left to check is that p-morphisms correspond to coalgebra morphisms. If $X,Y$ are Esakia spaces, and $p:X\to Y$ is a p-morphism, we need to show that:
$V_{G}(p)\circ\eta_{X}=\eta_{Y}\circ p.$
Note that these are two p-morphisms, and due to the universlity of the construction, two such p-morphisms will be equal if their projection on the first coordinate is the same; and indeed, we see that for $x\in X$:
$\pi_{0}\circ V_{G}(p)\circ\eta_{X}(x)=p(x)\text{ and }\pi_{0}\circ\eta_{Y}\circ p(x)=p(x).$
Which shows that the desired diagram commutes.
Conversely if $p:X\to Y$ is a coalgebra morphism, assume for $x\in X$ and $y\in Y$ that $p(x)\leq y$. Then $\eta_{Y}(p(x))\leq p(y)$, so $V_{G}(p)(\eta_{X}(x))\leq p(y)$. Hence by arguments similar to Lemma 3.12, we can find $x\leq x^{\prime}$ such that $p(x^{\prime})=y$, using the fact that we can find such an $x^{\prime}$ for finitely many coordinates. Hence $p$ is a p-morphism, as desired. ∎